For research purposes a sonic buoy is tethered to
Answer #1

V=(frequency)(wavelenght)
Where wavelength=V/frequency
= V/(1/T)
T=60.8 ms=60.8*10^-3 s= [60.8 / 1000] s

=V/(1/60.8*10^-3)

=VT=V*60.8*10^-3

=1522*60.8/1000=92.53 m

the wavelength of sound is 92.5 m